# Advection of the image point in probabilistically-reconstructed phase spaces

### v1

@@ Abstract
The HP method has been tested on the sea surface temperature and surface relative vorticity computed with the global 1/4 and 1/12 resolution NEMO model. Our results show that the HP solution (the solution computed with the HP method) in the probabilistically-reconstructed and reduced (in terms of dimensionality) phase space at 1/4 resolution is more accurate compared to the 1/4-solution computed with NEMO and it is also several orders of magnitude faster to compute than the 1/4-run. The proposed method shows encouraging results for the NEMO model and the potential for the use in other operational ocean and ocean-atmospheric models for both deterministic and probabilistic predictions.
@@
The importance of data-driven modelling in science and engineering is hard to underestimate. However, its applicability can be significantly limited due to the problem of data-deficiency. This problem can lead to inaccurate, low-fidelity results, or simply to inapplicability of the data-driven approach. The proposed study challenges the data-deficiency problem by offering to sample from the joint probability distribution of reference data to enrich the original reference data with new data from the reference data distribution. The method proposed in this work allows to harvest extra reference data to enable data-driven methods to work with meager datasets on which they would be inapplicable otherwise.
@@ 1 Introduction
The data-driven paradigm is gradually gaining momentum in computational fluid dynamics by offering an alternative to the physics-driven approach in situations when there is no physics-based model for the studied phenomenon or it is too computationally-expansive to acquire data from the model (e.g. (Mendez et al., 2023) and references therein). Being a powerful tool providing a wide spectrum of methods (e.g. Schneider et al. (2017); Zhang and Lin (2018); O’Gorman and Dwyer (2018); Zanna and Bolton (2020)) ranging from purely-data to physics-data driven, also know as hybrids, the data-driven approach typically suffers from the lack of reference data that can significantly compromise its fidelity or make it inapplicable. In order to help this problem, we offer a probabilistic reconstruction method which augments the hyper-parameterisation (HP) approach (e.g., Shevchenko and Berloff (2021, 2022a, 2022b, 2023a); Shevchenko and Crisan (2024)) with some ideas underlying the probabilistic-evolutionary approach (Shevchenko and Berloff, 2023b). More specifically, we propose to use Advection of the image point (a purely data-driven HP method working in phase space, also called state space) together with data sampled from the joint probability distribution of reference data so as to probabilistically reconstruct the reference phase space for the HP method and make it operable for short reference data records otherwise falling out of its range of applicability. Although, we consider the use of probabilistically-reconstructed phase spaces within the HP framework, its application is not bound to HP. For example, probabilistically-reconstructed reference datasets can be used within the context of other data-driven methods, as well as for training artificial neural networks, and for reconstruction of missing parts in observational data from different sources (drifters, weather stations, satellites, etc.)
@@ 2 The probabilistic reconstruction method: schematic and basics
(a) Reference data acquisition. The first step is to provide the method with reference data. Depending on the problem to solve, it can be numerical simulations, observations, or both. In our case it is sea surface temperature (SST) computed with a global 1/12 resolution ocean model NEMO and then interpolated onto a 1/4 grid (Fig. 1a). If observations are available then they can also be included in the reference dataset by simply adding them to the reference data record. In order to be consistent in space, observations must be on the same grid as the NEMO data; consistency in time is not required, as the proposed method does not use it.
@@
(b) Reference phase space calculation. The second step is to get the reference phase space for the HP method called ”Advection of the image point” (Shevchenko and Berloff, 2021, 2023a); the reference phase space is shown as an orange blob in Fig. 1b. The latter needs locations (which is SST itself) and directional vectors (SST tendencies), which are calculated from the reference data; different schemes can be used, we use the central finite difference in time. The tendencies are shown as black vectors in Fig. 1b, and the points they are attached to represent SST. The white spots (which we call voids) are regions of the phase space which lack data (usually, due to short reference records or unreliable observations). Typically, these voids is a reason for why data-driven methods cease to work on meager phase spaces.
@@
(c) Joint Probability Distribution calculation. The third step is to calculate the joint probability distribution (JPD) from the reference data in phase space; it is shown as a blue surface in Fig. 1c for illustration purposes; in high-dimensional phase spaces (like the one used for SST computed with NEMO) the JPD is a hyper-surface. It is important to remark that we calculate the JPD only for SST, while SST tendencies are computed in a different way explained below. The JPD for the reference SST is calculated globally with the histogram method, i.e. the whole reference dataset contributes to the histogram calculation. As the proposed method is developed to work in high-dimensional phase spaces, computing a multidimensional JPD and keeping it in memory for further sampling is an unaffordable option. Instead, we calculate a coordinate-wise probability density function (PDF). It gives us access to all necessary information for sampling at virtually no cost; we refer the reader to (Shevchenko and Berloff, 2023b) for a detail discussion on the JPD calculation. The only difference in the JPD calculation in this study is that we compute the JPD globally (i.e. for all reference data), while in (Shevchenko and Berloff, 2023b) it is computed locally in a given neighborhood.
(d) Sampling from JPD and calculation of probabilistically-reconstructed phase space. In the fourth step we sample SST from the JPD and then calculate SST tendencies by averaging reference SST tendencies over the neighborhood of sampled SST (Fig. 2). The localized calculation of tendencies allows to decrease the dimensionality of the JPD two times (as the JPD dimension for both SST and its tendencies is two times larger compared to the JPD only for SST) and thus make its calculation more efficient without compromising on accuracy. It is worth mentioning that tendencies can also be computed in a probabilistic way from a local JPD which is constructed based on the information from the neighborhood of the new tendency (red contour in Fig. 2), but this method is slower than the one used in this study.
@@
In order to sample from the JPD we use the inverse transform sampling method (Devroye, 1986). We have also tried the rejection sampling but did not gain that much of a difference in its favour; detailed explanations on sampling from the JPD are given in (Shevchenko and Berloff, 2023b). Sampling from the JPD allows to probabilistically reconstruct the originally sparsely-populated reference phase space by filling it with new locations and directional vectors (tendencies) needed to advect the image point (Fig. 1d). Note that the probabilistically-reconstructed reference data can be used not only in the HP method, but in other data-driven methods which suffer lack of data.
@@
It is important to note that SST and its tendencies in the probabilistically-reconstructed phase space are not ordered in time. This is the moment when we need Advection of the image point to compute a trajectory (solution), i.e to order those points in time.
@@
(e) Advection of the image point in probabilistically-reconstructed phase space. In the fifth step, the probabilistically-reconstructed phase space is used by Advection of the image point to calculate the trajectory (red line in Fig. 1e). For the reader’s convenience, we briefly describe how Advection of the image point works. This method falls into the category of data-driven methods from the hyper-parametersation (HP) class which takes advantage of working in phase spaces, as opposed to the conventional methods operating in the physical space. The method has been tested on both idealized and comprehensive ocean models (two-layer quasi-geostrophic model in a channel and MITgcm in the North Atlantic configuration) and showed significant improvements of the HP solution toward the reference one (Shevchenko and Berloff, 2021, 2023a). The HP approach currently umbrellas five different methods (ranging from purely data-driven to hybrids, which combine the data- and physics-driven paradigms). Another striking feature of the HP approach is that its measure of goodness is how close the HP solution is to the reference phase space. To put it another way, the HP approach matches phase spaces (the reference phase space and the phase space where the HP solution evolves), whereas conventional methods match individual trajectories. This measure allows the HP solution to evolve in the neighborhood of the reference phase space, and therefore reproduce the flow dynamics which is very similar to the reference one.
@@
$\frac{d\mathbf{y}}{dt}=\frac{1}{M}\sum\limits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{y}(t))}+\eta\left(\frac{1}{N}\sum\limits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{y}(t))}-\mathbf{y}(t)\right),\quad\mathbf{y}(t_{0})=\mathbf{y}_{0},$ (1)
@@
In a nutshell, the actual dynamics is given by the observed reference tendency $\frac{1}{M}\sum\nolimits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{y}(t))}$ the imperfect reconstruction of which from the reference data is compensated by nudging towards the observed reference state $\frac{1}{N}\sum\nolimits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{y}(t))}$ in the probabilistically-reconstructed phase space. Note that the neighbourhood is computed as the average over $M$ (and $N$ for the nudging term) nearest, in $l_{2}$ norm, to the solution $\mathbf{y}(t)$ points. The choice of the norm and the way the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is calculated is not limited to those used in this study, and can vary depending on what is needed. Our choice is, probably, the simplest, but it serves well the purpose of this work. We dub the set of parameters $\{M,N,\eta\}$ hyper-parameters. The hyper-parameters can be set based on the chosen metric and available data, we will return to their choice later. The interested reader is referred to (Shevchenko and Berloff, 2021, 2023a) for a more detailed discussion of the method.
@@
(f) Evolution in physical space. The fifth step is to get the modelled solution back in the physical space (Fig. 1f). In this study, the solution computed with the HP method is reshaped from a vector (used in the phase space) into a matrix, which is used to present fields in the physical space.
@@ 3.1 Chua’s circuit
As a minimum working example, we consider Chua’s circuit and demonstrate in more detail how the method works. Chua’s circuit (Chua et al., 1986) is given by the following system of equations:
@@
with $\mathbf{x}(t)=(x(t),y(t),z(t))$, and $\alpha=10$, $\beta=15$, $\gamma=0.01$, $a=0.1$, $c=-0.2$. As an initial condition, we take $\mathbf{x}(t_{0})=(1,0,0)$ so that the solution is close to the attractor.
Reference data acquisition. Within the context of Chua’s circuit, the reference data for the HP method is the solution $\mathbf{x}(t)$ and its tendencies $\mathbf{F}(\mathbf{x}(t))$. In order to get the reference data we integrate the Chua system (3) over time $t\in[0,100]$ and save the solution. Then, the tendencies are computed by differentiating $\mathbf{x}(t)$; we use the central finite difference in time (Fig. 4). Instead of differentiating the solution, one can use the right hand side, but we recommend against doing that, as the right hand side is usually unavailable for numerical simulations and observations. Note that instead of or in combination with the numerical solution, one can also use observations. In this case, one should form a new dataset $\widehat{\mathbf{x}}=\{\mathbf{x}(t),\mathbf{x}_{o}(t)\}$, which includes both the solution $\mathbf{x}(t)$ and observations $\mathbf{x}_{o}(t)$.
@@
The next step is to sample from the JPD to probabilistically reconstruct the reference phase space for the HP method. We use the same amount of reference data as before, but harvest twice as much from the JPD (Figs. 6,7). The results clearly demonstrate the high quality of the probabilistically-reconstructed data, and that it covers a larger phase space compared with the reference one.
@@
Figure 7: The same as in Fig. 6, but for the tendencies. Note that the tendencies are not sampled directly from the JPD but are computed locally in the neighbourhood of the sampled solution.

### v2

@@ Abstract
The HP method has been tested on the sea surface temperature and surface relative vorticity computed with the global 1/4 and 1/12 resolution NEMO model. Our results show that the HP solution (the solution computed with the HP method) in the probabilistically-reconstructed and reduced (in terms of dimensionality) phase space at 1/4 resolution is more accurate compared to the 1/4-solution computed with NEMO and it is also several orders of magnitude faster to compute than the 1/4-run. The proposed method shows encouraging results for the NEMO model and the potential for the use in other operational ocean and ocean-atmospheric models for both deterministic and probabilistic predictions. The method can also be used as a fast reanalysis in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution, as well as a dynamic interpolation method to stitch gaps in observational data.
keywords:
hyper-parameterization , NEMO ocean model , data-driven method , probabilistic reconstruction , phase space , state space , reanalysis , dynamic interpolation
journal: Journal of Computational Physics
@@
organization=National Oceanography Centre,addressline=European Way, city=Southampton, postcode=SO14 3ZH, country=UK
@@ 1 Introduction
The data-driven paradigm is gradually gaining momentum in computational fluid dynamics by offering an alternative to the physics-driven approach in situations when there is no physics-based model for the studied phenomenon or it is too computationally-expensive to acquire data from the model (e.g. [1] and references therein). The data-driven approach includes a variety of methods, spreading across different fields (e.g. statistical modelling [2, 3], dynamical system reconstruction [4, 5, 6, 7], hybrid modelling [8, 9], probabilistic-evolutionary approach [10], etc.) In recent years, machine learning (ML) methods started to enrich the data-driven approach with artificial neural networks (e.g. [11, 12, 13, 14, 15, 16]) typically trained on reanalysis data or reference runs of physics-driven models. The majority of these ML methods are used for predictive modelling (e.g. [17, 18, 19]), while some, known as Interpretable ML, go beyond traditional ML predictions and try to improve our understanding of modelled phenomena (e.g. [20, 21]).
@@
Being a powerful tool providing a wide spectrum of methods ranging from purely-data to physics-data driven, also know as hybrids, the data-driven approach typically suffers from the lack of reference data that can significantly compromise its fidelity or make it inapplicable. In order to help this problem, we offer a probabilistic reconstruction method which augments the hyper-parameterisation (HP) approach (e.g., [22, 8, 23, 24, 9]) with some ideas underlying the probabilistic-evolutionary approach [10]. More specifically, we propose to use Advection of the image point (a purely data-driven HP method working in phase space, also called state space) together with data sampled from the joint probability distribution of reference data so as to probabilistically reconstruct the reference phase space for the HP method and make it operable for short reference data records otherwise falling out of its range of applicability. The main difference between the proposed methodology and ML methods (based on artificial neural networks) is that: the former does not have the training phase (the calibration procedure of the HP method might be interpreted as such though), it uses phase spaces instead of the physical space, respects the regional stability property, and relies on the unconventional calculation of directional vectors. The regional stability is defined as a property of the solution of a differential equation to permanently remain in a neighborhood of the phase space region if the initial condition is sufficiently close to that region. By the unconventional calculation of directional vectors we mean how the right hand side (RHS) in the HP method is calculated. Namely, we do not use a sequence of following each other states from the past (as conventional integrators) to calculate the RHS, but use the states neighboring the current state in the phase space.
The main advantages of the HP method in probabilistically-reconstructed phase spaces are the following:
$\bullet$ Flexibility. The method enables research situations involving HP of only some part/component of the ocean circulation: geographical part (ocean surface or deep layers, different basins, etc.), specific field (salinity, temperature, velocity, etc.), specific observations (drifter, satellites, etc.), or their combination. For example, one can use HP only for the ocean circulation and solve separately for the biochemical processes embedded into it (e.g. carbon cycle research);
$\bullet$ Scalability. HP offers a great resolution scalability and can be used to parametrise the effect of any high-resolution process into lower-resolution systems (e.g. mesoscales processes in larger-scale processes, sub-mesoscale processes in mesoscale-resolving/sub-mesoscale-permitting systems, etc.), and therefore it can be readily adopted by forecasting systems;
$\bullet$ Alternative to parameterisations, which can be used in comprehensive ocean models (NEMO, MITgcm, FESOM, etc.);
$\bullet$ Fast generation of ensembles. HP is well-suited for generating ensembles of solutions, which serve many purposes in ocean and atmospheric modelling;
$\bullet$ HP as a tool for process studies. HP allows one to study oceanic feedbacks on climate. In this case the ocean is fully under control; one can systematically change parameters and properties of the ocean circulation and find out their global consequences. Using HP one can uncover and understand internal links and feedbacks within the system, by replacing specific components with HP analogues.
Although, we consider the use of probabilistically-reconstructed phase spaces within the HP framework, its application is not bound to HP. For example, probabilistically-reconstructed reference datasets can be used within the context of other data-driven methods, as well as for training artificial neural networks, and for reconstruction of missing parts in observational data from different sources (drifters, weather stations, satellites, etc.)
@@ 2 The probabilistic reconstruction method: schematic and basics
(a) Reference data acquisition. The first step is to provide the method with reference data. Depending on the problem to solve, it can be numerical simulations, observations, or both. In our case it is sea surface temperature (SST) computed with a global 1/12 resolution ocean model NEMO and then interpolated onto a 1/4 grid (Fig. 1a); SST is a scalar field defined on a cartesian grid. If observations are available then they can also be included in the reference dataset by simply adding them to the reference data record. In order to be consistent in space, observations must be on the same grid as the NEMO data; consistency in time is not required, as the proposed method does not use it.
(b) Reference phase space calculation. Before proceeding with the second step it would be helpful to remind the reader that a phase space (or state space) is an abstract mathematical space of all possible states of a physical system, while the physical space is typically defined as the actual three-dimensional space we all live in.
@@
The second step is to get the reference phase space for the HP method called ”Advection of the image point” [22, 24]; the reference phase space is shown as an orange blob in Fig. 1b. The latter needs states (which is SST itself) and directional vectors (time derivatives of SST, also referred to as tendencies), which are calculated from the reference data; different schemes can be used, we use the central finite difference in time. The tendencies are shown as black vectors in Fig. 1b, and the points they are attached to represent SST, which is a vector in the phase space, while it is a scalar field in the physical space.
The white spots (which we call voids) are regions of the phase space which lack data (usually, due to short reference records or unreliable observations). Typically, these voids is a reason for why data-driven methods cease to work on meager phase spaces.
@@
(c) Joint Probability Distribution calculation. The third step is to calculate the joint probability distribution (JPD) from the reference data in phase space; it is shown as a blue surface in Fig. 1c for illustration purposes. In high-dimensional phase spaces (like the one used for SST) the JPD is a hyper-surface. It is important to remark that we calculate the JPD only for SST, while SST tendencies are computed in a different way explained below. The JPD for the reference SST is calculated globally, i.e. the whole reference dataset is used in the calculation.
The standard approach to calculate an approximation to the JPD is to use the histogram method. This method is based on the subdivision of the domain into hypercubes (bins) and count the number of states in every bin to compute the height of each column of the multidimensional histogram; the domain is a hypercube in the phase space of SST in our case, the phase space dimension is equal to the number of grid points needed to approximate SST in the physical space. This method works well only in low-dimensional spaces, since the computer memory needed grows as the number of bins to the power of the space dimensions. Therefore, in high-dimensional spaces (like the one used for SST) computing a multidimensional histogram and keeping it in memory for further sampling is an unaffordable option.
In this study, we do not calculate the whole multi-dimensional histogram. Instead, we compute a coordinate-wise probability density function (PDF) for every component of the state (i.e. for every grid point of SST) and then use these PDFs to sample from the JPD; it gives us access to all necessary information for sampling at virtually no cost. Notice that these PDFs can be interpolated if there is not enough reference data to accurately approximate the PDFs.
(d) Sampling from JPD and calculation of probabilistically-reconstructed phase space. In the fourth step we sample SST from the JPD and then calculate SST tendencies by averaging reference SST tendencies over the neighborhood of sampled SST (Fig. 2).
@@
The localized calculation of tendencies allows to decrease the dimensionality of the JPD two times (as the JPD dimension for both SST and its tendencies is two times larger compared to the JPD only for SST) and thus make its calculation more efficient without compromising on accuracy. It is worth mentioning that tendencies can also be computed in a probabilistic way from a local JPD which is constructed based on the information from the neighborhood of the new tendency (red contour in Fig. 2), but this method is slower than the one used in this study.
@@
In order to sample from the JPD we use the inverse transform sampling method [25]. We have also tried the rejection sampling but did not gain that much of a difference in its favour; a sampling method based on a reverse-time stochastic differential equation [26] might be another option instead of the inverse transform sampling. Sampling from the JPD allows us to probabilistically reconstruct the originally sparsely-populated reference phase space by filling it with new states and directional vectors (tendencies) needed to advect the image point (Fig. 1d). Note that the probabilistically-reconstructed reference data can be used not only in the HP method, but in other data-driven methods which suffer lack of data.
We would like to draw the reader’s attention to the fact that the new states generated by sampling from the JPD does not mean new reference states (i.e. the states generated by the underlying reference model, it is the NEMO model in our case), since sampling from the JPD cannot guarantee that sampled states are reference states, and the HP dynamics in probabilistically-reconstructed phase spaces cannot guaranteed it either.
For a more detailed discussion on how the sampling procedure calculates a new state, we consider a two-dimensional case (i.e. a set of points in an XY-plane) for simplicity. In this case we have two coordinate-wise PDFs (one for each dimension). Sampling starts from computing the PDF for the first dimension and its corresponding CDF. Next, we draw a random number $r_{1}\in U[0,1]$, with $U[0,1]$ being a uniform distribution in the unit interval, and find $x_{1}(r_{1})$ with the inverse transform sampling method, where $x_{1}(r_{1})$ is the first coordinate of the new state. In order to calculate the second coordinate of the new state, we compute the PDF for the second dimension given $x_{1}(r_{1})$ and its CDF. Then, we draw another random $r_{2}\in U[0,1]$ and compute $x_{2}(r_{2})$, which is the second coordinate of the new state. In this way we can sample as many states as we need from the reference JPD.
@@
It is also important to note that SST and its tendencies in the probabilistically-reconstructed phase space are not ordered in time. This is the moment when we need Advection of the image point to compute a trajectory (solution), i.e to order those points in time.
@@
(e) Advection of the image point in probabilistically-reconstructed phase space. In the fifth step, the probabilistically-reconstructed phase space is used by Advection of the image point to calculate the trajectory (red line in Fig. 1e). For the reader’s convenience, we briefly describe how Advection of the image point works. This method falls into the category of data-driven methods from the hyper-parametersation (HP) class which takes advantage of working in phase spaces, as opposed to the conventional methods operating in the physical space. The method has been tested on both idealized and comprehensive ocean models (two-layer quasi-geostrophic model in a channel and MITgcm in the North Atlantic configuration) and showed significant improvements of the HP solution toward the reference one [22, 24]. The HP approach currently umbrellas five different methods (ranging from purely data-driven to hybrids, which combine the data- and physics-driven paradigms). Another striking feature of the HP approach is that its measure of goodness is how close the HP solution is to the reference phase space. To put it another way, the HP approach matches phase spaces (the reference phase space and the phase space where the HP solution evolves), whereas conventional methods match individual trajectories. This measure allows the HP solution to evolve in the neighborhood of the reference phase space, and therefore reproduce the flow dynamics which is very similar to the reference one.
@@
$\frac{d\mathbf{y}}{dt}=\frac{1}{M}\sum\limits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{y}(t)\right),\quad\mathbf{y}(t_{0})=\mathbf{y}_{0},$ (1)
@@
In a nutshell, the actual dynamics is given by the observed reference tendency $\frac{1}{M}\sum\nolimits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})$ the imperfect reconstruction of which from the reference data is compensated by nudging towards the observed reference state $\frac{1}{N}\sum\nolimits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}$ in the probabilistically-reconstructed phase space. Note that the neighbourhood is computed as the average over $M$ (and $N$ for the nudging term) nearest, in $l_{2}$ norm, to the solution $\mathbf{y}(t)$ points. The choice of the norm and the way the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is calculated is not limited to those used in this study, and can vary depending on what is needed. Our choice is, probably, the simplest, but it serves well the purpose of this work. We dub the set of parameters $\{M,N,\eta\}$ hyper-parameters. The hyper-parameters can be set based on the chosen metric and available data, we will return to their choice later. The interested reader is referred to [22, 24] for a more detailed discussion of the method.
@@
It is important to differentiate between the reference data set and the reference phase space. The former is the skeleton of the latter, i.e. the reference data set is a set of states used in the proposed methodology to discover the reference phase space through the use of the HP method in the probabilistically-reconstructed phase space.
@@
(f) Evolution in physical space. The fifth step is to compute the modelled solution back in the physical space (Fig. 1f). In this study, the solution computed with the HP method is reshaped from a vector (used in the phase space) into a scalar field, which is used to present fields in the physical space.
@@ 3.1 Chua’s circuit
As a minimum working example, we consider Chua’s circuit and demonstrate in more detail how the method works. Chua’s circuit [27] is given by the following system of equations:
@@
with $\mathbf{x}(t)=(x(t),y(t),z(t))$, and $\alpha=10$, $\beta=15$, $\gamma=0.01$, $a=0.1$, $c=-0.2$. As an initial condition, we take $\mathbf{x}(t_{0})=(1,0,0)$ so that the solution is close to the attractor. Note that the solution $\mathbf{x}(t)$ at time $t$ (being voltages across the capacitors and the current through the coil) is in the physical space, while the whole evolution of the solution (together with the time derivatives of the solution) is in the phase space.
Reference data acquisition. Within the context of Chua’s circuit, the reference data for the HP method is the solution $\mathbf{x}(t)$ and its tendencies $\mathbf{F}(\mathbf{x}(t))$. In order to get the reference data we integrate the Chua system (3) over time $t\in[0,100]$ and save the solution.
@@
Then, the tendencies are computed by differentiating $\mathbf{x}(t)$; we use the central finite difference in time (Fig. 4). Instead of differentiating the solution, one can use the right hand side, but we recommend against doing that, as the right hand side is usually unavailable for numerical simulations and observations. Note that instead of or in combination with the numerical solution, one can also use observations. In this case, one should form a new dataset $\widehat{\mathbf{x}}=\{\mathbf{x}(t),\mathbf{x}_{o}(t)\}$, which includes both the solution $\mathbf{x}(t)$ and observations $\mathbf{x}_{o}(t)$.
@@
The next step is to sample from the JPD to probabilistically reconstruct the reference phase space for the HP method. We use the same amount of reference data as before, but harvest twice as much from the JPD (Figs. 6,7).
@@
Figure 7: The same as in Fig. 6, but for the tendencies. The tendencies are not sampled directly from the JPD but are computed locally in the neighbourhood of the sampled solution.
The results clearly demonstrate the high quality of the probabilistically-reconstructed data, and that it covers a larger phase space compared with the reference one.